# Vision LLMs Are Bad at Hierarchical Visual Understanding, and LLMs Are the Bottleneck

### v1

@@ 3.1 Language Prompts Are Not the Bottleneck
Prompt engineering often comes as a remedy for boosting VLLMs’ performance in different applications [6, 55, 68, 58]. Could it also rescue VLLMs on our hierarchical visual understanding tasks? We strive to test prompt variants comprehensively. We specify the taxonomy levels in the prompts for CUB-200 [54] and iNat21 [53], whose taxonomies are grounded in biology. We even add CUB-200’s complete taxonomy as a JSON file to the prompts. For the other datasets with more generic taxonomies, we test general and chain-of-thought [24, 57] prompts derived from the template in Section 2.3. Appendix C provides all prompts in detail, and Figure 2 shows the results of some high-performing prompts. We can see from the results that the prompt design alone is insufficient to improve VLLMs’ hierarchical consistency or leaf-level accuracy.
@@ 3.2 Visual Embeddings Are Not the Bottleneck
The open-source VLLMs in this work vary in specific implementations, but their core components are the same: A visual encoder mapping images to embeddings, a projector translating visual embeddings into the language token space, and an LLM. If the hierarchical structure and discriminativeness are lost before the visual embeddings reach LLMs, the overall VLLMs would inevitably perform poorly on our hierarchical visual understanding tasks. Hence, it is crucial to examine the visual embeddings. We train three linear classifiers per taxonomy level to respectively probe the visual encoder, projector, and last layer of an LLM, where the image representations are an average of the visual tokens. Further details and results of the probing are provided in Appendix C.
Figure 3 shows the probing results of Qwen2.5-VL-7B over CUB-200 [54] and iNat21-Plant [53]. Remarkably, the linear classifiers outperform Qwen2.5-VL-7B all around. They achieve not only higher leaf-level accuracy than Qwen2.5-VL but also much better hierarchical consistency, even though the classifiers of different taxonomy levels are independently trained. Moreover, the linear probing results remain about the same at different stages of the forward propagation (i.e., immediately after the visual encoder, projector, and last layer of the VLLM), indicating that the visual tokens remain discriminative and structurally rich throughout different LLM layers. These results are a strong defense for the visual embeddings: They carry sufficient hierarchical and discriminative cues and should not be blamed for VLLMs’ poor hierarchical visual understanding performance.
3.3 LLMs Are the Bottleneck in VLLMs’ Hierarchical Visual Understanding
The huge discrepancy between the results of linearly probing visual tokens and VLLM performance in Figure 3 propels us to investigate other potential causes of VLLMs’ low hierarchical consistency beyond the visual embeddings, and we find that the influential LLMs are the bottleneck.
@@ 3.3 LLMs Are the Bottleneck in VLLMs’ Hierarchical Visual Understanding
[…]
@@
3.3.1 Open-Source VLLMs’ LLMs Lack Taxonomy Knowledge
We separate LLMs from open-source VLLMs and examine how much they know about the taxonomies used in our experiments. Mechanically, we reformulate our VQA tasks to a text-only version by replacing the images with their corresponding leaf labels:
Given the <leaf node label> (e.g., Anemone Fish), what is its taxonomicclassification at the <hierarchy> (e.g., kingdom) level?A.<similar class>  B.<ground truth>  C.<similar class>  D.<similar class>Answer with the option letter only.(Choices are shuffled in the $\begin{array}[]{l}\text{{Given the <leaf node label> (e.g., Anemone Fish), what is its taxonomic}}\\ \text{{classification at the <hierarchy> (e.g., kingdom) level?}}\\ \text{{A.<similar class> \ B.<ground truth> \ C.<similar class> \ D.<similar class>}}\\ \text{{Answer with the option letter only.}}\quad\text{(Choices are shuffled in the experiments)}\end{array}$
This process results in about 0.7 million QA tasks after deduplication. We use them to assess LLMs and report the (text) HCA results in Table 3 — we use (text/visual) HCA to refer to LLMs/VLLMs’ performance on text/visual QA tasks for clarity. We find that Qwen2.5-VL-7B’s LLM achieves only 63.86% (text) HCA on CUB-200, whose taxonomy comprises merely four levels. The LLMs of LLaVA-OV and InternVL-2.5 give rise to even lower (text) HCAs on CUB-200 (33% and 49%). One might wonder if these low (text) HCAs are due to that the biology taxonomy underlying CUB-200 is too specific for general LLMs.
@@ 3.3.1 Open-Source VLLMs’ LLMs Lack Taxonomy Knowledge
However, Table 3 further reveals that the LLMs also cannot perform well on ImgNet’s general taxonomies. Besides, we progressively simplify our QA tasks by chopping the iNat21-Plant taxonomy level by level. Figure 4 plots the (text) HCA results, which increase as the taxonomy becomes shallower (and, correspondingly, the leaf nodes are less fine-grained). Still, they are below 90% regardless of the taxonomies’ depths. There are noticeable drops at Levels 3 and 5 for Qwen2.5-VL and LLaVA-OV’s LLMs, implying that they pose more challenges than the other levels for the LLMs’ hierarchical reasoning. These results are surprising to a large degree, given the recent success of LLMs over various benchmarks and domains [1, 50, 60, 33, 45].
Correlation between (text) HCA and $\mathrm{Acc}_{\mathrm{leaf}}$-scaled (visual) HCA. An LLM’s low (text) HCA undoubtedly discounts its corresponding VLLM’s hierarchical consistency on visual inputs. We can quantify this notion using Pearson’s correlation coefficient. Since the (text) HCA’s corresponding leaf-level accuracy is 100% — we replaced images with their ground-truth leaf labels when making the text QA tasks, we normalize (visual) HCA by $1/\mathrm{Acc}_{\mathrm{leaf}}$. The last column in Table 3 shows that the correlation between (text) HCA and $\mathrm{Acc}_{\mathrm{leaf}}$-scaled (visual) HCA is as high as 0.9116.
A note about GPT-4o’s (text) HCA. The analyses above apply to only open-source VLLMs because we cannot separate LLMs from the proprietary GPT-4o. Unlike the open-source LLMs’ low (text) HCA, GPT-4o’s (text) HCA scores are as high as 98.81. Hence, the LLM part is not GPT-4o’s bottleneck in hierarchical visual understanding; instead, there are other possible causes of GPT-4o’s hierarchical inconsistency about the visual world.
3.3.2 Why Are LLMs Poor at Hierarchical Text Classification?
@@
Figure 5: Left: (Text) HCA difference between vision-language-tuned LLMs and the original ones. Right: (Text) HCA of linearly probing different layers of Qwen-2.5-VL-7B’s LLM on iNat21-Plant.
We answer this question by studying the original LLMs from which VLLMs are initialized, using the same text-only hierarchical classification setup described in Section 3.3. Figure 5 (Left) compares LLaVA-OV-7B and Qwen2.5-VL-7B’s LLMs with their corresponding original LLMs. First of all, we see that the original LLMs are on par with or even worse than their vision-tuned counterparts, indicates that the standalone LLMs still lack a strong grasp of taxonomy knowledge. Interestingly, Qwen2.5-VL’s LLM actually outperforms its original LLM on all taxonomies; in other words, the vision-language tuning actually enhances the LLM’s (text) hierarchical consistency. In contrast, LLaVA-OV’s vision-language tuning weakens the LLM’s (text) HCA.
LLMs Encode Hierarchical Structures Effectively but Cannot Decode Them Sufficiently. Next, we shift attention to the LLM embeddings of the concepts in our taxonomies — if the embeddings do not provide sufficient hierarchical structural cues, there is little chance LLMs can decode them. To this end, we convert a taxonomy into language prompts of three variants:
Prompt 1: <leaf node label> (e.g., Blue Jay) belongs to the <hierarchy>(e.g., Order) of <ground truth> (e.g., Passeriformes).Prompt 2: Given the <leaf node label>, what is its taxonomic classificationat the <hierarchy> level? It belongs to <ground truth>. Prompt 3: Given the <leaf node label>, what is its taxonomic classificationat the <hierarchy> $\begin{array}[]{l}\text{{Prompt 1:} {<leaf node label> (e.g., Blue Jay) belongs to the <hierarchy>}}\\ \text{{(e.g., Order) of <ground truth> (e.g., Passeriformes).}}\\ \text{{Prompt 2:} {Given the <leaf node label>, what is its taxonomic classification}}\\ \text{{at the <hierarchy> level? It belongs to <ground truth>. }}\\ \text{{Prompt 3:} {Given the <leaf node label>, what is its taxonomic classification}}\\ \text{{at the <hierarchy> level?}}\\ \end{array}$
We then train a linear classifier for each taxonomy level to probe the average embedding of the language tokens in every layer of an LLM. Figure 5 (Right) summarizes the (text) HCA results of Qwen2.5-VL-7B’s LLM on iNat21-Plant: The text embeddings give rise to highly hierarchically consistent linear probes. Especially for Prompt 3, with the ground-truth hierarchy labels withheld, the linear probes that receive only the leaf node embeddings can still achieve near-perfect hierarchical consistency in the LLM’s deeper layers. In other words, the specialized linear probes can decode the taxonomy knowledge significantly better than the general-purpose LLM.
LLMs’ Hierarchical Orthogonality Does Not Guarantee Hierarchical Consistency. Park et al. [42] recently predicted that LLMs represent hierarchical relations orthogonally in the representation space, e.g., animal is orthogonal to bird--mammal. They validated the prediction using Gemma [51] and LLaMA [19], and we further verify it in Figure 6 using both the original Qwen2.5-7B and the one after vision-language tuning. This pleasant geometric interpretation is, unfortunately, shadowed by the poor performance of Gemma and Qwen2.5-7B on our taxonomy QA tasks — we report the Gemma results in Appendix C. We argue that more fine-grained analyses of the LLM representation are required to establish a relationship between LLMs’ hierarchical consistency and geometry.
@@
4 LLMs Gain More Hierarchical Consistency than VLLMs from Finetuning
Could we improve the VLLMs’ hierarchical visual understanding capabilities via finetuning using our VQA tasks built upon taxonomies? Likely, no, because LLMs are the bottleneck: The LLMs’ hierarchical consistency over text-only tasks is so bad (Table 3) that we conjecture this shortcoming can only be fixed in the pretraining stage rather than the “tail patching” finetuning stage.
Still, the following presents some LoRA-finetuning [22] experiments with Qwen2.5-VL-7B, the best-performing 7B VLLM in our previous experiments, mainly for two reasons. One is to see how much finetuning could help, even though we believe pretraining instead of finetuning should be the rescue to VLLMs’ hierarchical inconsistency. The other is further to investigate the interplay between VLLMs and their LLMs — interestingly, our results reaffirm that LLMs are the bottleneck for VLLMs’ hierarchical visual understanding because LLMs’ performance gain from the finetuning upper-bounds VLLMs’. Our finetuning data consists of VQA tasks constructed from iNat21-Plant’s training set, covering 3,771 species nodes in the taxonomy instead of the full 4,271 species nodes. We then evaluate the finetuned model’s improvement on iNat21-Plant, its generalization to other hierarchical visual understanding datasets, and how well it maintains the general vision-language capabilities. Please see Appendix D for more details on the training.
@@ 4 LLMs Gain More Hierarchical Consistency than VLLMs from Finetuning
Table 5: (Text) HCA of the LLM of Qwen2.5-VL-7B before and after the LoRA-finetuning.
Model iNat21-Animal iNat21-Plant ImgNet-Animal CUB-200
LLM of Qwen2.5-VL-7B 52.08 64.21 68.14 63.86
LLM of Qwen2.5-VL-7B (LoRA) 65.63 84.87 72.39 66.15
@@
Results and Discussion. Tables 4 shows that finetuning Qwen2.5-VL using the VQA tasks that partially cover the iNat21-Plant taxonomy delivers improvements on both iNat21-Plant and other datasets. On iNat21-Plant, HCA rises from $17.67$ to $29.34$ ($+11.67$ absolute gain), while $\mathrm{Acc_{leaf}}$ gains $6.05$. The HCA on ImageNet-Animal increases from $56.00$ to $58.62$ and on CUB-200 from $43.76$ to $46.17$. More interestingly, Table 5 indicates that the LLM’s (text) HCA increases more from the finetuning than Qwen2.5-VL’s (visual) HCA (e.g., $20.66$ vs. $11.67$ on iNat21-Plant and $4.25$ vs. $2.62$ on ImgNet-Animal). To some extent, this finding reaffirms that LLMs are the bottleneck of VLLMs’ hierarchical visual understanding, and one has to improve LLMs’ (text) taxonomy knowledge to boost VLLMs’ (visual) hierarchical consistency. Besides, our results demonstrate that vision-language training can benefit both VLLMs and their LLMs, aligning with some recent advocates for improving LLMs using multimodal data beyond language only [31, 52]. Appendix D reports more results and discussion, including that the finetuned model does not lose its general capability tested on MME [16], MMBench [36], and SEED-Bench [30].
5 Related Work
Hierarchical classification [47, 25] enables many applications. It is vital for a comprehensive understanding of the visual world [61, 43, 65, 48, 7, 44] and many language concepts [70, 56, 71, 21]. Several recent studies have revisited this longstanding problem and shown that CLIP-style [46] models lack consistency across taxonomic levels [58, 18]. Wu et al. [58] evaluate CLIP under multiple levels of semantic granularity and introduce a hierarchy-consistent prompt tuning method. Pal et al. [40] enhance CLIP’s hierarchical representations by embedding them to a hyperbolic space. Xia et al. [59] further extend this direction by incorporating graph-based representation learning. Novack et al. [38] use hierarchical information to improve zero-shot classification accuracy. Zhang et al. [68] first identified the limitations of current VLLMs in fine-grained image classification. Building on this, Liu et al. [35] further assess a broader range of VLLMs. He et al. [20] point out a potential cause, the scarcity of image class names in pretraining. Beyond closed-set evaluation [63, 17], Conti et al. [12] benchmark VLLMs’ open-world classification, while Snæbjarnarson et al. [49] propose to evaluate VLLMs’ open-set predictions using a taxonomic similarity rather than exact string matching. However, to the best of our knowledge, no prior work has examined VLLMs under the hierarchical visual understanding context.

### v2

@@ 4.1 Language Prompts Are Not the Bottleneck
Prompt engineering often comes as a remedy for boosting VLLMs’ performance in different applications [8, 65, 80, 68]. Could it also rescue VLLMs on our hierarchical visual recognition tasks? We strive to test prompt variants comprehensively. We specify the taxonomy levels in the prompts for CUB-200 [64] and iNat21 [63], whose taxonomies are grounded in biology. We even add CUB-200’s complete taxonomy as a JSON file to the prompts. For the other datasets with more generic taxonomies, we test general and chain-of-thought [28, 67] prompts derived from the template in Section 3.3. Appendix C provides all prompts in detail, and Fig.˜2 shows the results of some high-performing prompts. We can see from the results that prompt design alone is insufficient to improve VLLMs’ hierarchical consistency or leaf-level accuracy.
@@ 4.2 Visual Embeddings Are Not the Bottleneck
LLM of iNat21-Animal iNat21-Plant ImgNet-Artifact ImgNet-Animal CUB-200 $\rho(\text{text,visual})$
@@
The open-source VLLMs in this work vary in specific implementations, but their core components are the same: A visual encoder mapping images to embeddings, a projector translating visual embeddings into the language token space, and an LLM. If the hierarchical structure and discriminativeness are lost before the visual embeddings reach LLMs, the overall VLLMs would inevitably perform poorly on our hierarchical visual recognition tasks. Hence, it is crucial to examine the visual embeddings. We train three linear classifiers per taxonomy level to respectively probe the visual encoder, projector, and last layer of an LLM, where the image representations are an average of the visual tokens. See Appendix C for further details and results.
Figure 3 shows the probing results of Qwen2.5-VL-7B over CUB-200 and iNat21-Plant. Remarkably, the linear classifiers outperform Qwen2.5-VL-7B all around. They achieve not only higher leaf-level accuracy than Qwen2.5-VL but also much better hierarchical consistency, even though the classifiers of different taxonomy levels are independently trained. Moreover, the linear probing results remain about the same at different stages of the forward propagation (i.e., immediately after the visual encoder, projector, and last layer of the VLLM), indicating that the visual tokens remain discriminative and structurally rich throughout different LLM layers. These results are a strong defense for the visual embeddings: They carry sufficient hierarchical and discriminative cues and should not be blamed for VLLMs’ poor hierarchical visual recognition performance.
4.3 LLMs Are the Bottleneck in Open-Source VLLMs’ Hierarchical Visual Recognition
The huge discrepancy between the results of linearly probing visual tokens and VLLM performance in Fig.˜3 propels us to investigate other potential causes of VLLMs’ low hierarchical consistency beyond the visual embeddings, and we find that the LLMs are the bottleneck.
4.3.1 Open-Source VLLMs’ LLMs Lack Sufficient Taxonomy Knowledge
Text-Only Hierarchical Consistency. We separate LLMs from open-source VLLMs and examine how much they know about the taxonomies used in our experiments. Mechanically, we reformulate our VQA tasks to a text-only version by replacing the images with their corresponding leaf labels:
[…]
This process produces 0.7 million QA tasks after deduplication. We use them to assess LLMs’ text-only hierarchical consistency over our vision taxonomies.
@@ 4.3.1 Open-Source VLLMs’ LLMs Lack Sufficient Taxonomy Knowledge
Results. Table 2 shows the results, where we use (text/visual) HCA to refer to LLMs/VLLMs’ performance on text/visual QA tasks. We find that Qwen2.5-VL’s LLM achieves only 63.86% (text) HCA on CUB-200, whose taxonomy comprises merely four levels. The LLMs of LLaVA-OV and InternVL-2.5 give rise to even lower (text) HCAs on CUB-200 (33% and 49%).
Figure 5: Left: (Text) HCA difference between vision-language-tuned LLMs and original ones. Right: (Text) HCA of linearly probing different layers of Qwen-2.5-VL-7B’s LLM on iNat21-Plant.
One might wonder if those low (text) HCAs are due to that the biology taxonomy underlying CUB-200 is too specific for general LLMs. However, Table 2 further reveals that the LLMs also cannot perform well on ImgNet’s general taxonomies. Besides, we progressively simplify our QA tasks by chopping the iNat21-Plant taxonomy level by level. Figure 4 plots the (text) HCA results, which increase as the taxonomy becomes shallower (and, correspondingly, the leaf nodes are less fine-grained). Still, they are below 90% regardless of the taxonomies’ depths. There are noticeable drops at Levels 3 and 5 for Qwen2.5-VL and LLaVA-OV’s LLMs, implying that they pose more challenges than the other levels for the LLMs’ hierarchical reasoning. These results are highly unexpected, given the recent success of LLMs over various benchmarks and domains [47, 60, 72, 38, 71].
(text) HCA vs. (visual) HCA. An LLM’s low (text) HCA undoubtedly discounts its corresponding VLLM’s hierarchical consistency on visual inputs. We can quantify this notion using Pearson’s correlation coefficient. Since the (text) HCA’s corresponding leaf-level accuracy is 100% — we replaced images with their ground-truth leaf labels when making the text QA tasks, we normalize (visual) HCA by $1/\mathrm{Acc}_{\mathrm{leaf}}$. The last column in Tab.˜2 shows that the correlation between (text) HCA and $\mathrm{Acc}_{\mathrm{leaf}}$-scaled (visual) HCA is as high as 0.9116.
A note about GPT-4o’s (text) HCA. The analyses above apply to only open-source VLLMs because we cannot separate LLMs from the proprietary GPT-4o. Unlike the open-source LLMs’ low (text) HCA, GPT-4o’s (text) HCA scores are as high as 98.81. Hence, we conjecture that the LLM part is likely not GPT-4o’s bottleneck in hierarchical visual recognition; instead, there are other possible causes of GPT-4o’s hierarchical inconsistency about the visual world.
4.3.2 Open-Source LLMs Lack Taxonomy Knowledge
@@
We answer this question by studying the original LLMs from which VLLMs are initialized, using the same text-only hierarchical classification setup described in Section 4.3.1. Figure 5 (Left) compares InternVL2.5-8B and Qwen2.5-VL-7B’s LLMs with their corresponding original LLMs. First of all, we see that the original LLMs are on par with or even worse than their vision-tuned counterparts, indicating that the standalone LLMs still lack a strong grasp of taxonomy knowledge. Interestingly, Qwen2.5-VL’s LLM outperforms its original LLM on every taxonomy, and InternVL2.5-8B’s LLM also improves on four of five; in other words, the vision-language tuning actually enhances the LLM’s (text) hierarchical consistency.
LLMs Encode Hierarchical Structures Effectively but Cannot Decode Them Sufficiently. Next, we shift attention to the LLM embeddings of the concepts in our taxonomies — if the embeddings do not provide sufficient hierarchical structural cues, there is little chance LLMs can decode them. To this end, we convert a taxonomy into language prompts of four variants:
[…]
We then train a linear classifier for each taxonomy level to probe the average embedding of the language tokens in every layer of an LLM. Figure˜5 (Right) summarizes the (text) HCA results of Qwen2.5-VL-7B’s LLM on iNat21-Plant: The text embeddings give rise to highly hierarchically consistent linear probes. Even for Prompt 3 (ground-truth hierarchy labels withheld) and Prompt 4 (hierarchy-free input), the linear probes that receive only the leaf node embeddings can still achieve near-perfect hierarchical consistency in the LLM’s deeper layers. In other words, the specialized linear probes can decode the taxonomy knowledge significantly better than the general-purpose LLM.
LLMs’ Hierarchical Orthogonality Does Not Guarantee Hierarchical Consistency. Park et al. [51] recently predicted that LLMs represent hierarchical relations orthogonally in the representation space, e.g., animal is orthogonal to bird-mammal. They validated the prediction using Gemma [61] and LLaMA [21], and we further verify it in Fig.˜6 using both the original Qwen2.5-7B and the one after vision-language tuning. This pleasant geometric interpretation is, unfortunately, shadowed by the poor performance of Gemma and Qwen2.5-7B on our taxonomy QA tasks — we report the Gemma results in Appendix C. We argue that more fine-grained analyses of the LLM representation are required to establish a relationship between LLMs’ hierarchical consistency and geometry.
5 Visual Finetuning Cannot Fix LLMs
Could we improve the VLLMs’ hierarchical visual recognition capabilities via finetuning using our VQA tasks built upon taxonomies? Likely, no, because LLMs are the bottleneck: The LLMs’ hierarchical consistency over text-only tasks is so bad (Tab.˜2) that we conjecture this issue can only be fixed in the pretraining stage rather than the “tail patching” finetuning stage, echoing the recent result in [23].
Still, the following presents some LoRA-finetuning [27] experiments with Qwen2.5-VL-7B, the best-performing 7B VLLM in our previous experiments, mainly for two reasons. One is to see how much finetuning could help, even though we believe pretraining instead of finetuning should be the rescue to VLLMs’ hierarchical inconsistency. The other is further to investigate the interplay between VLLMs and their LLMs — interestingly, our results reaffirm that LLMs are the bottleneck for VLLMs’ hierarchical visual recognition because LLMs’ performance gain from the finetuning upper-bounds VLLMs’. Our finetuning data consists of VQA tasks constructed from iNat21-Plant’s training set. We then evaluate the finetuned model’s improvement on iNat21-Plant, its generalization to other hierarchical visual recognition datasets, and how well it maintains the general vision-language capabilities. Please see Appendix D for more details on the training.
Results and Discussion. Table˜3 shows that finetuning Qwen2.5-VL using the VQA tasks that partially cover the iNat21-Plant taxonomy delivers improvements on both iNat21-Plant and other datasets. On iNat21-Plant, HCA rises from $17.67$ to $29.34$ (+11.67+11.67 absolute gain), while $\mathrm{Acc_{leaf}}$ gains $6.05$. The HCA on ImgNet-Animal increases from $56.00$ to $58.62$ and on CUB-200 from $43.76$ to $46.17$. More interestingly, Table˜4 indicates that the LLM’s (text) HCA increases more from the finetuning than Qwen2.5-VL’s (visual) HCA (e.g., $20.66$ vs. $11.67$ on iNat21-Plant and $4.25$ vs. $2.62$ on ImgNet-Animal). This finding reaffirms that LLMs are the bottleneck and one has to improve LLMs’ (text) taxonomy knowledge to boost VLLMs’ (visual) hierarchical consistency to some extent. Besides, our results demonstrate that vision-language training can benefit both VLLMs and their LLMs, aligning with some recent advocates for improving LLMs using multimodal data beyond language only [35, 62]. Appendix D reports more results, including that the tuned model does not lose its general capability tested on MME [18], MMBench [41], and SEED-Bench [34].
Table 4: (Text) HCA of the LLM of Qwen2.5-VL-7B before and after the LoRA fine-tuning.
LLM of iNat-Animal iNat-Plant ImgNet-Animal CUB-200
Qwen2.5-VL-7B 52.08 64.21 68.14 63.86
Qwen2.5-VL-7B (LoRA) 65.63 84.87 72.39 66.15